OpenStudy (anonymous):

(3/16+9/16)(1/4-5/9)

OpenStudy (anonymous):

-432/25 or -17.28.

OpenStudy (anonymous):

\[({3\over 16}+{9\over 16})({1\over 4}-{5 \over 9})\] \[({3+9\over 16})({1\cdot 9\over 4 \cdot 9}-{5\cdot 4 \over 9\cdot 4 })\] \[({12\over 16})({9\over 36}-{20 \over 36})\] \[({3\cdot 4\over 4\cdot 4})({9-20\over 36})\] \[({3\over 4})({-11\over 36})\] \[{3\cdot -11}\over{4\cdot 36}\] \[{3\cdot -11}\over{4\cdot 12 \cdot 3}\] \[{-11}\over{4\cdot 12}\] \[-11\over 48 \]

OpenStudy (anonymous):

-11/48